# A Deconfounding Approach to Climate Model Bias Correction

### v1

@@ Abstract
Global Climate Models (GCMs) are crucial for predicting future climate changes by simulating the Earth systems. However, GCM outputs exhibit systematic biases due to model uncertainties, parameterization simplifications, and inadequate representation of complex climate phenomena. Traditional bias correction methods, which rely on historical observation data and statistical techniques, often neglect unobserved confounders, leading to biased results. This paper proposes a novel bias correction approach to utilize both GCM and observational data to learn a factor model that captures multi-cause latent confounders. Inspired by recent advances in causality based time series deconfounding, our method first constructs a factor model to learn latent confounders from historical data and then applies them to enhance the bias correction process using advanced time series forecasting models. The experimental results demonstrate significant improvements in the accuracy of precipitation outputs. By addressing unobserved confounders, our approach offers a robust and theoretically grounded solution for climate model bias correction.
Introduction
Global Climate Models (GCMs), such as those developed by the Coupled Model Intercomparison Project phase 6 (CMIP6) (Eyring et al. 2016), are vital tools for predicting future climate changes. These models simulate the physical and chemical processes of the Earth systems—including the atmosphere, oceans, land, and ice—to provide detailed climate forecasts. Despite significant advancements in GCM, their output still exhibits systematic biases. These biases primarily come from uncertainties within the models, simplifications in parameterization processes, and inadequate representations of complex climate phenomena (Mouatadid et al. 2023). For instance, processes such as cloud formation, precipitation, radiation, and convection are often simplified into parameterization formulas in GCMs, which may not accurately reflect real-world conditions, thus leading to biases. Consequently, (Lafferty and Sriver 2023) highlighted that bias corrections were particularly crucial for near-term precipitation projections, especially in cases where observational data are inconsistent with GCMs output.
To enhance the reliability and accuracy of GCM outputs, numerous bias correction techniques, ranging from simple linear scaling to advanced quantile mapping, have been proposed to modify GCM outputs and align them more closely with actual observations (Casanueva et al. 2020; Chen et al. 2013; Dowdy 2020; Enayati et al. 2020; Feigenwinter et al. 2018; Heo et al. 2019; Lafon et al. 2013; Maraun 2016; Mehrotra, Johnson, and Sharma 2018; Miao et al. 2016; Nahar, Johnson, and Sharma 2018; Piani, Haerter, and Coppola 2010; Smitha et al. 2018; Teutschbein and Seibert 2012a; Wu et al. 2022b). These methods typically rely on historical observation data and statistical approaches to adjust climate model outputs and correct systematic errors (Maraun and Widmann 2018).
@@ Introduction
$\left[t-h-w-1,\ldots,t-w\right]$ $\left[t,t+1,\ldots,t+k+1\right]$
@@
Table 1: A description of data variables for GCM and observations.
@@
However, these traditional methods have been found to inflate simulated extremes, raising concerns about their use in climate change applications where extremes are significant, such as drought and flooding (Huang, Hall, and Berg 2014; Pastén-Zapata et al. 2020). Another major limitation of most bias correct methods is their assumption that all relevant factors are known and observable, which is unrealistic in practical applications. Climate systems are complex and include many factors that may not be fully observed, such as microclimate effects, regional climate characteristics, and anthropogenic influences. These potential unobserved (confounding) factors dynamically impact time series forecast. However, since they cannot be fully observed, they are frequently overlooked by current bias correction methods. This overlook limits the potential of most existing methods in bias correction.
@@
To deal with the challenge of extreme conditions, the study by (Nivron et al. 2024) incorporates advanced time series forecasting models into the bias correction of extreme weather events, such as heatwaves, by considering the bias correction as a time-indexed regression model with stochastic output. This provides a new perspective: adapting time series forecasting models to bias correction can improve model performance. However, although advanced models have shown significant potential in time series forecasting (Zhou et al. 2021; Wu et al. 2022a, 2023a; Wang et al. 2024), particularly in climate forecasting (Bi et al. 2023; Wu et al. 2023b), they typically overlook the existence of unobserved confounding factors, which leads to biased results.
@@
In the field of causal inference with time series, recent studies have started addressing the challenge of unobserved confounders in predicting the potential outcomes of treatments in time series data. Instead of assuming unconfoundedness (Pearl 2000), these studies operate under a weaker assumption that only multi-cause confounders exist (Bica, Alaa, and van der Schaar 2020; Wang and Blei 2019; Li et al. 2024; Cheng et al. 2023; Xu et al. 2023). A variable is said to be a confounder if it is a common cause of both the treatment and the outcome (Pearl 2009). For example, in the Directed Acyclic Graph (DAG) shown in Figure 1(a), the confounder $\mathbf{X}$ affects both the treatment $\mathbf{A}$ and outcome $Y$. In our bias correction problem, precipitation output is always treated as the outcome, and other related climate variables are treated as treatment variables. The most common method for predicting potential outcomes is to use all treatment variables. However, neglecting unobserved multi-cause confounders $\mathbf{Z}$, such as large-scale atmospheric circulation patterns or oceanic processes, can lead to simultaneous effects on various climate variables, including precipitation. These multi-cause confounders make it possible to apply current research outcomes to bias correction in GCM.
@@
Inspired by the work of (Bica, Alaa, and van der Schaar 2020), in this paper, we propose a deconfounding bias correction method. It is important to note that estimating hidden confounders in climate bias correction is much more complex. This increased complexity arises not only from the intricate nature of hidden confounders in climate science but also because these confounders need to be inferred from GCM and observation data.
@@
    We use causality to identify and understand unobserved confounders, allowing us to obtain unbiased outcome in the presence of multi-cause confounders. By identifying latent confounders through constructing a factor model over time, our approach can capture these unobserved factors, resulting in unbiased outcomes.
@@
    We develope a two-phase algorithm: Deconfounding and Correction. The Deconfounding Bias Correction (BC) factor model captures the confounders from both Global Climate Model (GCM) output data and observational data, and the correction model uses these confounders as additional information for the bias correction task. Inspired by the TSD (Time Series Deconfounder) (Bica, Alaa, and van der Schaar 2020), our method extend to GCM and observation data.
Problem Formulation
As shown in Table 1, let the random variable $\mathbf{V}$ represent the set of climate variables, such as humidity, temperature, pressure, precipitation and others. Among these variables, $Y$ denote the outcome of interest, for example, precipitation. In our problem setting, because we will use both historical and current data of these variables, to avoid confusion, $\mathbf{X}_{t}$ represent historical version of $\mathbf{V}\setminus{Y}$. $\mathbf{A}_{t}$ respresnt current version of $\mathbf{V}\setminus{Y}$.
@@
The data for a location, also known as the location trajectory, consist of realizations of the previously described random variables $\{\mathbf{x_{t}^{G}},\mathbf{a_{t}^{G}},y_{t}^{G},\mathbf{x_{t}^{O}},\mathbf{a_{t}^{O}},y_{t}^{O}\}$. Let $y(\mathbf{\bar{a}_{t}})$ represent the potential outcome (precipitation), which could be either factual or counterfactual, for each possible treatment course $\mathbf{\bar{a}_{t}}$, where $\mathbf{\bar{a}_{t}}=(\mathbf{a_{1}},...,\mathbf{a_{t}})$. Consequently, we have $y^{G}(\mathbf{\bar{a}_{t}})$ and $y^{O}(\mathbf{\bar{a}_{t}})$. The concept of potential outcomes allows us to consider what the outcome $y$ would be under different treatment scenarios, which is essential for causal inference (Rubin 1974).
@@
To correct future $k$ steps precipitation bias $\Delta{Y}=Y^{O}-Y^{G}$ which is the difference between observational precipitation $y^{O}$ and GCM output precipitation $y^{G}$, a common way is to use all data to obtain a regression model (Nivron et al. 2024) which can be represented by Figure 1(a). However, the existence of hidden confounders (e.g. unobserved atmospheric and oceanic circulation) will result in biased result. To address the issue of hidden confounders in time series data, (Bica, Alaa, and van der Schaar 2020) have developed a method, assuming the presence of multi-cause hidden confounders $\mathbf{Z}$, as shown in the Figure 1(b).
@@
$\mathbf{E}\left[\mathbf{Y}^{O}(\mathbf{\bar{a}}^{O}_{\geq t})\mid\mathbf{\bar{A}}^{O}_{t-1},\mathbf{\bar{X}}^{O}_{t},\mathbf{\bar{Z}}_{t}\right]$
Figure 2: Deconfounding BC factor model
Proposed method
The existence of confounders can result in the obtained association among $\mathbf{X,A},Y$ not accurately representing the true relationships, potentially leading to biased results. To solve this, we (1) input the climate model data into the Deconfounding Bias Correction (BC) factor model to obtain the multi-cause hidden confounder which is essential for bias correction (we call this step ’Deconfounding’). Then (2) we use a hidden multicause confounder as a bias source, combine it with observational data, building a precipitation correction model to help the climate model to have a better estimate of future output (we call this step ’Correction’).
Deconfounding
To address the challenge of deconfounding time series data with time varing latent confounder, (Bica, Alaa, and van der Schaar 2020) proposed the Time Series Deconfounder (TSD), which extends the deconfounder methodology introduced by (Wang and Blei 2019) to the time series domain. The fundamental principle of the TSD is that It utilizes factor rmodel to infer substitutes for hidden confounders as treatments.
The goal of this part is to generalize the factor model proposed by (Bica, Alaa, and van der Schaar 2020) to a Deconfounding BC factor model for bias correction. This extension aims to extend the Time Series Deconfounder to a complex time series setting of two sources.
Figure 3: Double source, $i$ denote as the total number of climate variables.
Deconfounding BC factor model
To extend time series deconfounder (Bica, Alaa, and van der Schaar 2020) to bias correction, We propose an enhanced multi-source factor model specially for this task, termed the Deconfounding BC factor model.
@@
$p(a_{t1},\ldots,a_{tk}\mid\mathbf{z}_{t},\mathbf{x}_{t})=\prod_{j=1}^{k}p(a_{tj}\mid\mathbf{z}_{t},\mathbf{x}_{t}).$
To extend this model for multi-source data, consider both GCM and Observations. The unobserved confounder will affect $\mathbf{X,A,Y}$ in both sources, allowing us to infer the sequence of unobserved confounders $\mathbf{z}_{t}=g(\bar{\mathbf{h}}_{t-1})$ that can be used to render both source treatments conditionally independent, where $\bar{\mathbf{h}}_{t-1}=\{{\bar{\mathbf{h}}_{t-1}^{O},\bar{\mathbf{h}}_{t-1}^{G}}\}$ is the realization of $\bar{\mathbf{H}}_{t-1}$. Specifically, the factorization for the multi-source scenario can be expressed as follows:
$p(a_{t1}^{O},\ldots,a_{tk}^{O}\mid\mathbf{z}_{t},\mathbf{x}_{t}^{O})=\prod_{j=1}^{k}p(a_{tj}^{O}\mid\mathbf{z}_{t},\mathbf{x}_{t}^{O}),$
$p(a_{t1}^{G},\ldots,a_{tk}^{G}\mid\mathbf{z}_{t},\mathbf{x}_{t}^{G})=\prod_{j=1}^{k}p(a_{tj}^{G}\mid\mathbf{z}_{t},\mathbf{x}_{t}^{G}).$
It allows us to infer the sequence of latent variables $\mathbf{Z}$ that can be used to render the treatments conditionally independent with the observed location covariates $\mathbf{x_{t}}$. As in one location, the same confounders $\mathbf{Z}$ wil affect both GCM and observations. In this case, the $\mathbf{Z}$ we learned should render treatments in both GCM and observational data.
Since the structure of the factor model depends on causality, which relies on the assumptions listed below.
@@
Positivity (Overlap): if $P(\mathbf{\bar{A}}_{t-1}=\mathbf{\bar{a}}_{t-1},\mathbf{X}_{t}=\mathbf{\bar{x}}_{t})\neq 0$ then $P(A_{t}=a_{t}\mid\mathbf{\bar{A}}_{t-1}=\mathbf{\bar{a}}_{t-1},\mathbf{X}_{t}=\mathbf{\bar{x}}_{t})>0$ for all $a_{t}$.
@@
$\mathbf{Y}(\bar{\mathbf{a}}_{\geq t})\ \perp\!\!\!\perp\ \mathbf{A}_{tj}\mid\mathbf{X}_{t},\bar{\mathbf{H}}_{t-1},\text{ for all }t\in\{0,\ldots,T\},\text{ and for all }j\in\{1,\ldots,k\}.$
Next, we provide a theoretical analysis for the soundness of the learned $\mathbf{Z}_{t}$ by introducing the concept of sequential Kallenberg construction (Bica, Alaa, and van der Schaar 2020), as follows.
@@
Sequential Kallenberg construction At timestep $t$, we say that the distribution of assigned causes $(\mathbf{A}^{G}_{t1},\ldots,\mathbf{A}^{G}_{tk}),(\mathbf{A}^{O}_{t1},\ldots,\mathbf{A}^{O}_{tk})$ admits a sequential Kallenberg construction from the random variables $\mathbf{Z}_{t}=g(\mathbf{\bar{H}}^{O}_{t-1},\mathbf{\bar{H}}^{G}_{t-1})$ and $\mathbf{X}^{O}_{t},\mathbf{X}^{G}_{t}$ if there exist measurable functions $f_{tj}^{O}:\mathbf{Z^{O}_{t}}\times\mathbf{X^{O}_{t}}\times[0,1]\to A_{j}^{O}$ and $f_{tj}^{G}:\mathbf{Z^{G}_{t}}\times\mathbf{X^{G}_{t}}\times[0,1]\to A_{j}^{G}$ and random variables $U_{tj}^{O},U_{tj}^{G}\in[0,1]$, with $j=1,\ldots,k$ such that:
$A_{tj}^{O}=f_{tj}^{O}(\mathbf{Z_{t}},\mathbf{X^{O}_{t}},U_{tj}^{O}),$
$A_{tj}^{G}=f_{tj}^{G}(\mathbf{Z_{t}},\mathbf{X^{G}_{t}},U_{tj}^{G}),$
@@
$(U_{t1}^{O},\ldots U_{tk}^{O})\perp Y^{O}(\mathbf{\bar{a}}^{O}_{\geq t})\mid\mathbf{Z_{t}},\mathbf{X^{O}_{t}},\mathbf{\bar{H}_{t-1}}^{O},$
$(U_{t1}^{G},\ldots U_{tk}^{G})\perp Y^{G}(\mathbf{\bar{a}}^{G}_{\geq t})\mid\mathbf{Z_{t}},\mathbf{X^{G}_{t}},\mathbf{\bar{H}_{t-1}}^{G},$
@@
Theorem 1 shows that the latent variables $\mathbf{Z_{t}}$ are infered using the fact causes $(A_{t1}^{G},\ldots,A_{tk}^{G})$ and $(A_{t1}^{O},\ldots,A_{tk}^{O})$ are jointly independent given $\mathbf{Z_{t}}$ and $\mathbf{X_{t}}^{G}$, $\mathbf{X_{t}}^{O}$ as shown in Figure 3. The result means that, at each timestep, for both sources, the variable $\mathbf{\bar{X}}_{t}$, $\mathbf{\bar{Z}}_{t}$, $\mathbf{\bar{A}}_{t-1}$ contain all of the dependencies between the potential outcomes $\mathbf{Y}(\mathbf{\bar{a}}_{\geq t})$ and the assigned causes $\mathbf{A}_{t}$. See the Appendix for the full proof.

### v2

# Deconfounding Multi-Cause Latent Confounders: A Factor-Model Approach to Climate Model Bias Correction

@@ Abstract
Global Climate Models (GCMs) are crucial for predicting future climate changes by simulating the Earth systems. However, the GCM Outputs exhibit systematic biases due to model uncertainties, parameterization simplifications, and inadequate representation of complex climate phenomena. Traditional bias correction methods, which rely on historical observation data and statistical techniques, often neglect unobserved confounders, leading to biased results. This paper proposes a novel bias correction approach to utilize both GCM and observational data to learn a factor model that captures multi-cause latent confounders. Inspired by recent advances in causality based time series deconfounding, our method first constructs a factor model to learn latent confounders from historical data and then applies them to enhance the bias correction process using advanced time series forecasting models. The experimental results demonstrate significant improvements in the accuracy of precipitation outputs. By addressing unobserved confounders, our approach offers a robust and theoretically grounded solution for climate model bias correction.
1 Introduction
Global Climate Models (GCMs), such as those developed by the Coupled Model Intercomparison Project phase 6 (CMIP6) Eyring et al. (2016), are vital tools for predicting future climate changes. These models simulate the physical and chemical processes of the Earth systems—including the atmosphere, oceans, land, and ice—to provide detailed climate forecasts. Despite significant advancements in GCM, their output still exhibits systematic biases. These biases primarily come from uncertainties within the models, simplifications in parameterization processes, and inadequate representations of complex climate phenomena Mouatadid et al. (2023). For instance, processes such as cloud formation, precipitation, radiation, and convection are often simplified into parameterization formulas in GCMs, which may not accurately reflect real-world conditions, thus leading to biases. Consequently, Lafferty and Sriver (2023) highlighted that bias corrections were particularly crucial for near-term precipitation projections, especially in cases where observational data are inconsistent with the GCM Output.
To enhance the reliability and accuracy of the GCM Outputs, numerous bias correction techniques, ranging from simple linear scaling to advanced quantile mapping, have been proposed to modify the GCM Outputs and align them more closely with actual observations Casanueva et al. (2020); Chen et al. (2013); Dowdy (2020); Enayati et al. (2020); Feigenwinter et al. (2018); Lafon et al. (2013); Maraun (2016); Mehrotra et al. (2018); Miao et al. (2016); Nahar et al. (2018); Piani et al. (2010); Smitha et al. (2018); Teutschbein and Seibert (2012); Wu et al. (2022b). These methods typically rely on historical observation data and statistical approaches to adjust climate model outputs and correct systematic errors Maraun and Widmann (2018).
@@ 1 Introduction
$\left[t-h-w-1,\ldots,t-w-1\right]$ $\left[t-w,\ldots,t\right]$ $\left[t+1,\ldots,t+k+1\right]$
@@
Table 1: A description of data variables for GCM and observations, and $\mathbf{V}=\{\text{humidity},\text{temperature},\text{pressure},\text{precipitation},\ldots\}$
@@
However, these traditional methods have been found to inflate simulated extremes, raising concerns about their use in climate change applications where extremes are significant, such as drought and flooding Pastén-Zapata et al. (2020). Another major limitation of most bias correct methods is their assumption that all relevant factors are known and observable, which is unrealistic in practical applications. Climate systems are complex and include many factors that may not be fully observed, such as microclimate effects, regional climate characteristics, and anthropogenic influences. These potential unobserved (confounding) factors dynamically impact time series forecast. However, since they cannot be fully observed, they are frequently overlooked by current bias correction methods. This overlook limits the potential of most existing methods in bias correction.
@@
To deal with the challenge of extreme conditions, the study by Nivron et al. (2024) incorporates advanced time series forecasting models into the bias correction of extreme weather events, such as heatwaves, by considering the bias correction as a time-indexed regression model with stochastic output. This provides a new perspective: adapting time series forecasting models to bias correction can improve model performance. However, although advanced models have shown significant potential in time series forecasting Zhou et al. (2021); Wu et al. (2022a, 2023a); Wang et al. (2024); Gao et al. (2024), particularly in climate forecasting Bi et al. (2023); Wu et al. (2023b), they typically overlook the existence of unobserved confounding factors, which leads to biased results.
@@
In the field of causal inference with time series, recent studies have started addressing the challenge of unobserved confounders in predicting the potential outcomes of treatments in time series data. Instead of assuming unconfoundedness Pearl (2000), these studies operate under a weaker assumption that only multi-cause confounders exist Bica et al. (2020); Wang and Blei (2019); Li et al. (2024); Cheng et al. (2023). A variable is said to be a confounder if it is a common cause of both the treatment and the outcome Hernán and Robins (2020). For example, in the Directed Acyclic Graph (DAG) shown in Figure 1(a), the confounder $\mathbf{X}$ affects both the treatment $\mathbf{A}$ and outcome $Y$. In our bias correction problem, precipitation output is always treated as the outcome, and other related climate variables are treated as treatment variables. The most common method for predicting potential outcomes is to use all treatment variables. However, neglecting unobserved multi-cause confounders $\mathbf{Z}$, such as large-scale atmospheric circulation patterns or oceanic processes, can lead to simultaneous effects on various climate variables, including precipitation. These multi-cause confounders make it possible to apply current research outcomes to bias correction in GCM.
@@
Inspired by the work of Bica et al. (2020), in this paper, we propose a deconfounding bias correction method. It is important to note that estimating hidden confounders in climate bias correction is much more complex. This increased complexity arises not only from the intricate nature of hidden confounders in climate science but also because these confounders need to be inferred from GCM and observation data.
@@
    We use causality to identify and understand unobserved confounders, allowing us to obtain an unbiased outcome in the presence of multi-cause confounders. By identifying latent confounders through constructing a factor model over time, our approach can capture these unobserved factors, resulting in unbiased outcomes.
@@
    We develop a two-phase algorithm: Deconfounding and Correction. The Deconfounding Bias Correction (BC) factor model captures the confounders from both Global Climate Model (GCM) output data and observational data, and the correction model uses these confounders as additional information for the bias correction task. Inspired by the TSD (Time Series Deconfounder) Bica et al. (2020), our method extends to GCM and observation data.
2 Problem Formulation
As shown in Table 1, let the random variable $\mathbf{V}$ represent the set of climate variables, such as humidity, temperature, pressure, precipitation and others. Among these variables, $Y$ denote the outcome of interest, for example, precipitation. In our problem setting, because we will use both historical and current data of these variables, to avoid confusion, $\mathbf{X}_{t}$ represents the historical version of $\mathbf{V}\setminus{Y}$. $\mathbf{A}_{t}$ represents the current version of $\mathbf{V}\setminus{Y}$.
@@ 2 Problem Formulation
The data for a location, also known as the location trajectory, consist of realizations of the previously described random variables $\{\mathbf{x_{t}^{G}},\mathbf{a_{t}^{G}},y_{t}^{G},\mathbf{x_{t}^{O}},\mathbf{a_{t}^{O}},y_{t}^{O}\}$. Let $y(\mathbf{\bar{a}_{t}})$ represent the potential outcome (precipitation), which could be either factual or counterfactual, for each possible treatment course $\mathbf{\bar{a}_{t}}$, where $\mathbf{\bar{a}_{t}}=(\mathbf{a_{1}},...,\mathbf{a_{t}})$. Consequently, we have $y^{G}(\mathbf{\bar{a}_{t}})$ and $y^{O}(\mathbf{\bar{a}_{t}})$. The concept of potential outcomes allows us to consider what the outcome $y$ would be under different treatment scenarios, which is essential for causal inference Rubin (1974).
@@
To correct future $k$ steps precipitation bias $\Delta{Y}=Y^{O}-Y^{G}$ which is the difference between observational precipitation $y^{O}$ and GCM output precipitation $y^{G}$, a common way is to use all data to obtain a regression model Nivron et al. (2024) which can be represented by Figure 1(a). However, the existence of hidden confounders (e.g. unobserved atmospheric and oceanic circulation) will result in biased result. To address the issue of hidden confounders in time series data, Bica et al. (2020) have developed a method, assuming the presence of multi-cause hidden confounders $\mathbf{Z}$, as shown in the Figure 1(b).
@@
$\mathbf{E}\left[\mathbf{Y}^{O}(\mathbf{\bar{a}}^{O}_{\geq t})\mid\mathbf{\bar{A}}^{O}_{t-1},\mathbf{\bar{X}}^{O}_{t},\mathbf{\bar{Z}}_{t}\right]$ (1)
3 Proposed Method
The existence of confounders can result in the obtained association among $\mathbf{X,A},Y$ not accurately representing the true relationships, potentially leading to biased results. To solve this, we (1) input the climate model data into the Deconfounding Bias Correction (BC) factor model to obtain the multi-cause hidden confounder which is essential for bias correction (we call this step ’Deconfounding’). Then (2) we use a hidden multicause confounder as a bias source, combine it with observational data, building a precipitation correction model to help the climate model to have a better estimate of future output (we call this step ’Correction’) [^1] Code is avaliable at github: https://github.com/Wentao-Gao/A-Factor-Model-Approach-to-Climate-Model-Bias-Correction.
3.1 Deconfounding
To address the challenge of deconfounding time series data with time-varying latent confounders, Bica et al. (2020) proposed the Time Series Deconfounder (TSD), which extends the deconfounder methodology introduced by Wang and Blei (2019) to the time series domain. The fundamental principle of the TSD is that it utilizes factor model to infer substitutes for hidden confounders as treatments.
The goal of this part is to generalize the factor model proposed by Bica et al. (2020) to a Deconfounding BC factor model for bias correction. This extension aims to extend the Time Series Deconfounder to a complex time series setting of two sources.
Figure 2: Causal graphs of two sources with latent confounders.
3.1.1 Deconfounding BC Factor Model
To extend time series deconfounder Bica et al. (2020) to bias correction, We propose an enhanced multi-source factor model specially for this task, termed the Deconfounding BC factor model.
@@ 3.1.1 Deconfounding BC Factor Model
$p(a_{t1},\ldots,a_{tk}\mid\mathbf{z}_{t},\mathbf{x}_{t})=\prod_{j=1}^{k}p(a_{tj}\mid\mathbf{z}_{t},\mathbf{x}_{t}).$ (2)
Figure 3: The architecture of the proposed Deconfounding BC factor model. At each time step $t$, a shared RNN takes the past treatment and covariate sequences from both GCM and observational sources as input and generates a latent confounder representation $\mathbf{Z}_{t}$. This latent confounder, along with the current covariates from GCM ($\mathbf{X}_{t}^{G}$) and observations ($\mathbf{X}_{t}^{O}$), is then fed into multiple fully connected layers (one per treatment variable) to predict the treatment distributions $p(A_{tj}^{G}\mid\mathbf{X}_{t}^{G},\mathbf{Z}_{t})$ and $p(A_{tj}^{O}\mid\mathbf{X}_{t}^{O},\mathbf{Z}_{t})$. By structuring the model, the learned $\mathbf{Z}_{t}$ serves to deconfound both sources simultaneously, enabling conditionally independent treatment modeling given $\mathbf{Z}_{t}$ and observed covariates.
To extend this model for multi-source data, consider both GCM and Observations as in Figure 2. The unobserved confounder will affect $\mathbf{X,A,Y}$ in both sources, allowing us to infer the sequence of unobserved confounders $\mathbf{z}_{t}=g(\bar{\mathbf{h}}_{t-1})$ that can be used to render both source treatments conditionally independent, where $\bar{\mathbf{h}}_{t-1}=\{{\bar{\mathbf{h}}_{t-1}^{O},\bar{\mathbf{h}}_{t-1}^{G}}\}$ is the realization of $\bar{\mathbf{H}}_{t-1}$. Specifically, the factorization for the multi-source scenario can be expressed as follows:
$p(a_{t1}^{O},\dots,a_{tk}^{O}\mid\mathbf{z}_{t},\mathbf{x}_{t}^{O})=\prod_{j=1}^{k}p(a_{tj}^{O}\mid\mathbf{z}_{t},\mathbf{x}_{t}^{O}).$ (3)
$p(a_{t1}^{G},\ldots,a_{tk}^{G}\mid\mathbf{z}_{t},\mathbf{x}_{t}^{G})=\prod_{j=1}^{k}p(a_{tj}^{G}\mid\mathbf{z}_{t},\mathbf{x}_{t}^{G}).$ (4)
It allows us to infer the sequence of latent variables $\mathbf{Z}_{t}$ such that, conditioned on $\mathbf{Z}_{t}$ and the covariates from both sources $\mathbf{X}_{t}^{G}$ and $\mathbf{X}_{t}^{O}$, the treatments become conditionally independent. As in one location, the same confounders $\mathbf{Z}$ will affect both GCM and observations. In this case, the $\mathbf{Z}$ we learned should render treatments in both GCM and observational data conditionally independent with covariate $\mathbf{X_{t}}$.
The structure of the factor model depends on causality, which relies on the assumptions listed below.
@@
Positivity (Overlap)Bica et al. (2020): if $P(\mathbf{\bar{A}}_{t-1}=\mathbf{\bar{a}}_{t-1},\mathbf{X}_{t}=\mathbf{\bar{x}}_{t})\neq 0$ then $P(A_{t}=a_{t}\mid\mathbf{\bar{A}}_{t-1}=\mathbf{\bar{a}}_{t-1},\mathbf{X}_{t}=\mathbf{\bar{x}}_{t})>0$ for all $a_{t}$.
@@
$\mathbf{Y}(\bar{\mathbf{a}}_{\geq t})\perp\!\!\!\perp{A}_{tj}\mid\mathbf{X}_{t},\bar{\mathbf{H}}_{t-1}$ (5)
[…]
Next, we provide a theoretical analysis for the soundness of the learned $\mathbf{Z}_{t}$ by introducing the concept of sequential Kallenberg construction Bica et al. (2020), as follows.
@@
Sequential Kallenberg construction At timestep $t$, we say that the distribution of assigned causes $({A}^{G}_{t1},\ldots,{A}^{G}_{tk}),({A}^{O}_{t1},\ldots,{A}^{O}_{tk})$ admits a sequential Kallenberg construction from the random variables $\mathbf{Z}_{t}=g(\mathbf{\bar{H}}^{O}_{t-1},\mathbf{\bar{H}}^{G}_{t-1})$ and $\mathbf{X}^{O}_{t},\mathbf{X}^{G}_{t}$ if there exist measurable functions $f_{tj}^{O}:\mathbf{Z^{O}_{t}}\times\mathbf{X^{O}_{t}}\times[0,1]\to A_{j}^{O}$ and $f_{tj}^{G}:\mathbf{Z^{G}_{t}}\times\mathbf{X^{G}_{t}}\times[0,1]\to A_{j}^{G}$ and random variables $U_{tj}^{O},U_{tj}^{G}\in[0,1]$, with $j=1,\ldots,k$ such that:
$A_{tj}^{O}=f_{tj}^{O}(\mathbf{Z_{t}},\mathbf{X^{O}_{t}},U_{tj}^{O})$ (6)
$A_{tj}^{G}=f_{tj}^{G}(\mathbf{Z_{t}},\mathbf{X^{G}_{t}},U_{tj}^{G}),$ (7)
@@
$\bigl{(}U_{t1}^{O},\dots,U_{tk}^{O}\bigr{)}\;\mathrel{\perp\!\!\!\perp}\;Y^{O}\bigl{(}\bar{\mathbf{a}}^{O}_{\geq t}\bigr{)}\;\bigm{|}\;\mathbf{Z}_{t},\;\mathbf{X}_{t}^{O},\;\bar{\mathbf{H}}_{t-1}^{O}$ (8)
$\bigl{(}U_{t1}^{G},\dots,U_{tk}^{G}\bigr{)}\;\mathrel{\perp\!\!\!\perp}\;Y^{G}\bigl{(}\bar{\mathbf{a}}^{G}_{\geq t}\bigr{)}\;\bigm{|}\;\mathbf{Z}_{t},\;\mathbf{X}_{t}^{G},\;\bar{\mathbf{H}}_{t-1}^{G}$ (9)
@@
Theorem 1 establishes that the latent variable $\mathbf{Z}_{t}$ can be inferred by leveraging the conditional independence structure of the treatments: within each source (GCM or Observation), the assigned treatments $\{A_{t1},\ldots,A_{tk}\}$ are conditionally independent given $\mathbf{Z}_{t}$ and their respective covariates $\mathbf{X}_{t}^{G}$ and $\mathbf{X}_{t}^{O}$. This structure, illustrated in Figure 3, provides the necessary identifiability condition for inferring $\mathbf{Z}_{t}$ through a shared factor model across both sources. The result means that, at each timestep, for both sources, the variable $\mathbf{\bar{X}}_{t}$, $\mathbf{\bar{Z}}_{t}$, $\mathbf{\bar{A}}_{t-1}$ contain all of the dependencies between the potential outcomes $\mathbf{Y}(\mathbf{\bar{a}}_{\geq t})$ and the assigned causes $\mathbf{A}_{t}$.